PROBING THE MOLECULAR DYNAMICS OF A Cu(CD_{3}OD) CLUSTER WITH PHOTODETACHMENT-PHOTOIONIZATION SPECTROSCOPY

\underline{J. BARBERA}, V. DRIBINSKI, and W.C. LINEBERGER, JILA/Department of Chemistry, University of Colorado, Boulder, CO 80309; S. HORVATH, and A.B. McCOY, Department of Chemistry, Ohio State University, Columbus, OH 43210.

We report the femtosecond nuclear dynamics of neutral Cu(CD_{3}OD) van der Waals clusters, investigated using photodetachment-photoionization spectroscopy. Photodetachment of Cu(CD_{3}OD) anion with a 150 fs 400 nm laser pulse produces a vibrationally excited neutral complex that undergoes ligand reorientation and dissociation. The time evolving neutral is interrogated by delayed femtosecond resonant two photon ionization. This study shows that the nascent Cu(CD_{3}OD) complex dissociates on prompt (3 ps) and slower timescales (30 ps). The prompt component reflects direct dissociation upon photodetachment, while the slower dissociation arises from the coupling of CD_{3}OD molecular rotation into the Cu-(CD_{3}OD) dissociation coordinate. Theoretical investigations provide insight to the nature of the molecular dynamics which produce the observed dissociation characteristics. Supported by NSF and AFOSR